### v1

# Active Matter as a framework for living systems-inspired Robophysics

Abstract
Robophysics investigates the physical principles that govern living-like robots operating in complex, real-world environments. Despite remarkable technological advances, robots continue to face fundamental efficiency limitations. At the level of individual units, locomotion remains a challenge, while at the collective level, robot swarms struggle to achieve shared purpose, coordination, communication, and cost efficiency. This perspective article examines the key challenges faced by bio-inspired robotic collectives and highlights recent research efforts that incorporate principles from active-matter physics and biology into the modeling and design of robot swarms.
I Introduction
The development of robots has been a growing focus of the scientific community for over half a century, with particularly rapid advances in the past three decades Aguilar et al. (2016); Wang et al. (2021). These versatile, remotely controlled, or autonomous devices are increasingly essential due to their ability to perform tasks that are dangerous, repetitive, or difficult for humans. Robots offer significant benefits in tasks such as search and rescue operations Lyu et al. (2023), being able to access unsafe environments, as well as damage evaluation, agricultural monitoring, forest management Ecke et al. (2022), and crowd monitoring Husman et al. (2021); Conte et al. (2021); Al-Sheary and Almagbile (2017); Gohari et al. (2022), enhancing precision and efficiency. Despite their remarkable capabilities and diverse applications, robots still encounter substantial efficiency limitations both as individual units and in collective operations. In comparison to biological organisms, at the single-robot level, locomotion remains a critical challenge: robots often struggle to navigate natural terrains, such as dense forests, rubble-strewn areas, or regions with gusty windsAguilar et al. (2016); Schaeffer and Lindstedt (2013); Fajen (2013); Biewener and Patek (2018). Additionally, mission planning is frequently challenging due to the robot’s inability to effectively manage unforeseen obstacles, resulting in decreased operational efficiency and reliability. To address these challenges, recent research efforts have focused on incorporating principles from physics and biology into the modeling and design of robotsVolpe et al. (2025).
The gap in understanding motion within complex environments, particularly from a physics perspective, has given rise to the field of Robophysics Aguilar et al. (2016). Robophysics is defined as the pursuit of principles governing the movement and control of robot-like entities in intricate environments. It delves into fundamentally non-equilibrium processes where motion emerges through internal actuation Levine and Goldman (2023). Advancing the understanding of robotic movement principles necessitates the application of experimental physics, combining theoretical and numerical frameworks with systematically controlled laboratory experiments on simplified systems.
In this perspective article, we examine how Robophysics provides a framework for understanding and improving both individual and collective robotic behaviors. We first discuss how insights from living systems have informed the design of single-robot locomotion strategies. We then explore the extension of these principles to collective systems, highlighting the challenges that arise in robot swarms. Finally, we outline how recent advances in active matter physics and machine learning can be integrated to develop predictive, adaptive, and purpose-driven robotic collectives.
II Learning from living systems: single robot
At the single-robot level, Robophysics has focused on understanding and improving locomotion by building simplified devices capturing essential geometric and control features of biological systems Aguilar et al. (2016). These devices function as abstract models, providing insights into appendage-induced locomotion, blurring the dichotomy between failure and success to uncover principles of environment-locomotor interaction Aguilar et al. (2016). Consequently, the exploration of robotic motion serves as an exemplary model system for advancing our understanding of complex dynamical systems driven far from equilibrium Bechinger et al. (2016); Volpe et al. (2025). So far, the scientific literature has reported both terrestrial and aerial examples in which Robophysics played a pivotal role in enabling robots to adapt to real-world conditions by incorporating living-like mechanisms into their design.
On the one side, on land, snake-like robots have been designed to move through unpredictable and confined spaces by mimicking the undulating motion of snakes, which use their body segments and frictional forces to generate propulsion Takemori et al. (2018); Tanaka et al. (2015). This ability makes snake-like robots highly suitable for deployment in disaster zones, where they can traverse rubble and debris to locate survivors or assess structural damageQiyuan and Chen (2020); Fu et al. (2020). Studies of snake locomotion across diverse terrains have allowed the creation of robots excelling in mobility, adaptability, and durability, even under extreme conditions Moosavi et al. (2025); Thanabalan et al. (2025). At a smaller length-scale, peristaltic warm-inspired soft robots have been designed for a better understanding of locomotion in different environments Das et al. (2023).
On the other side, in the air, bird-inspired drones have been tailored to exploit flapping-wing aerodynamics to achieve maneuverability and energy efficiency which surpasses traditional fixed-wing or rotor designsPhan and Floreano (2024). At a smaller length scale, insect-inspired robots have been designed for crop pollination and environmental monitoringChen et al. (2019). Such bio-inspired abstractions demonstrate how Robophysics, by bridging biology and engineering, enhances robotic adaptability in complex real-world environments Jeger et al. (2024).
In what follows, we show that a similar approach can be taken at the collective level. Just as individual robots benefit from bio-inspired locomotion strategies, robot swarms can benefit from coordination and adaptability principles observed in animal groups and other living ensembles.
III Learning from living systems: collective behaviour
Living systems achieve collective functionality not via a central control, but via local interactions which generate robust and adaptable group behavior Hauser (1997); Gillam (2011). Across length-scales, organisms actively consume energy to remain out of equilibrium and thereby dynamically respond to environmental cues Wang et al. (2015). For example, cytoskeletal filaments continuously assemble and disassemble to enable dynamic cell shape changes Fletcher and Mullins (2010), while flocks of birds or schools of fish reorganize within fractions of a second to conserve energy or evade predators Marchetti et al. (2013); Couzin et al. (2005); Portugal et al. (2014). The unifying principle is that coordinated responses emerge from simple, local rules of sensing and interaction, yet giving rise to collective resilience and adaptability.
Efficient robot swarms rely on fundamental processes-communication, coordination, cooperation, collaboration, and, at times, competition Prorok et al. (2021), to achieve collective goals. Algorithms based on these tasks have already achieved remarkable results in controlled experimental environments with a limited number of individuals in the swarm (fewer than ten). The following sections discuss each of these aspects and their relevance to collective robotic behavior.
III.1 Communication and its constraints
Gielis et al. Gielis (2022) argue that communication in robot swarms confronts two major challenges. First, learning models should be employed not only to determine what information to communicate and to whom, but also to optimize multiple aspects of swarm performance simultaneously. Second, robots must be designed together with their communication capabilities, since any control algorithm relying on explicit communication is prone to failure when network data are insufficient and to sub-optimal performance when the network transmits excessive information.
At a first glance, Robophysics may appear similar to bio-inspired networking, as both draw on concepts from collective behavior in natural systems. However, bio-inspired networking primarily aims to leverage biological models to improve the performance of routing, coordination, synchronization, and resource-allocation algorithms in large-scale communication networks Dressler and Akan (2010). Robophysics, instead, seeks to develop a physics-based approach of how real robots move, interact, and generate collective behaviors under physical constraints. From a computer networking perspective, the operation of a swarm of small, low-cost robots faces several fundamental challenges. Unlike abstract network nodes, robots are physical entities whose positions and motions directly affect network connectivity. The network topology continuously evolves with the motion of the swarm, requiring tightly coupled network strategies for locomotion and control.
However, a location-deprived environment, as is typical in many indoor robophysical experiments, poses an additional challenge to robots in terms of simultaneous location and mapping (SLAM), which is necessary for each robot to estimate both its relative position and the overall network topology. Typical SLAM systems assume relatively high-cost, high-performance hardware (e.g., LiDAR) and often cannot be deployed effectively on ultra-low-cost robotic platforms due to high power, mechanical complexity, and computational demands.
Performing topology mapping and cooperative localization in indoor environments with only simple sensors is thus a significant challenge. Relying on low-cost sensors inherently introduces sensor uncertainty, since measurements of distance, velocity, or orientation are often noisy or incomplete. This, in turn, leads to perception uncertainty, as obstacles or neighboring robots may be only partially detected or misidentified due to occlusions or reflections. Moreover, robot mobility exacerbates communication uncertainty, as time-varying and intermittent wireless links can cause packet losses and transmission delays.
In Robophysics, however, communication is not merely an abstract tool but an integral component of the physical process itself: information exchange influences motion, and motion, in turn, shapes connectivity. Modeling and controlling this bidirectional feedback loop remains an open challenge situated at the intersection of network dynamics, control theory, and collective behavior.
III.2 Coordination and Scalability
Beyond communication, an equally critical challenge lies in coordination, which seeks to minimize redundant tasks and interference through cooperative behaviors such as path synchronization for area coverage and collision avoidance Yao et al. (2022).
Designing a robot swarm requires scalability as the number of robots increases. Yet, many techniques that perform well for small groups become impractical when the swarm size grows. Recent research has therefore focused on the formal control of shape formation in robot collectives Sakurama and Sugie (2021). In these systems, robots communicate with their local neighbors to maintain a desired configuration, and their connections are typically represented as a graph. Scaling up such formations involves enlarging the graph while preserving its topological properties, a task that remains particularly challenging.
Figure 1: Schematic illustration of how principles from living systems can inform robophysics. In living systems (top), groups of organisms such as fish, ants, and birds display collective behavior, adaptability, and shared purpose. These principles inspire robotic collectives (bottom), where multi-robot swarms must address challenges such as communication, perception, and sensing, and energy and resource management.
III.3 Cooperation, routing and planning
Cooperation refers to the ability of similar agents to work together on tasks that cannot be accomplished individually, such as jointly moving a payload or assisting one another in crossing a gap. For many years, the field of computer networking has drawn ideas from bio-inspired models of collective behavior, leading to the development of algorithmic paradigms such as ant-colony-inspired routing and optimization schemes. In ant-based routing protocols for mobile ad hoc networks, artificial ’ants’ (data packets) explore paths and probabilistically reinforce successful ones, mimicking pheromone trail deposition. This approach, known as Ant Colony Optimization (ACO), has proven effective in dynamic, decentralized settings by leveraging self-organization and adaptability to changing network topologies.
Similarly, evolutionary algorithms such as genetic algorithms (GAs) have been applied to path planning and task assignment in multi-drone systems. For example, GA-based strategies like GenPath Bartolini et al. (2021) optimize UAV (Unmanned Aerial Vehicle) missions under constraints such as limited battery life, intermittent communication, and multi-phase recharging, enabling scalable coordination in search-and-rescue, environmental monitoring, and parcel delivery.
While ant colony optimization and genetic algorithms excelled at solving static or well-defined problems of routing and planning, they did not capture the real-time, embodied responsiveness of biological collectives, where coordination emerges from local physical interactions and rapid adaptation to changing conditions.
This gap highlights the relevance of Robophysics to adopt a different point of view: learning from living systems to program robots whose collective behavior emerges in analogous ways. According to this view, coordination is not imposed top-down but arises from the living-inspired physics of local interactions, fluctuations, and energy consumption.
III.4 Collaboration, Learning and Adaptation
Collaboration emphasizes teamwork with a diverse team that leverages the unique skills of different individuals, such as integrating various sensors or carrying out a common patrol task combining aerial and aquatic autonomous vehicles.
Learning concerns the swarm’s ability to learn from new scenarios and adjust its collective behaviour to achieve its goals, which is crucial for effective operation. While learning, robots in a swarm can collectively perceive their surroundings and share that information among themselves (distributed perception and sensing).
To better adapt to the environment, robots should exhibit both robustness and fault tolerance. Although these properties are related, they represent distinct concepts. Robustness refers to the ability to maintain satisfactory performance even when environmental conditions differ from those anticipated during the design phase. Fault tolerance, instead, refers to the capacity of the swarm to sustain effective operation despite failures in individual agents, sensors, or other components.
III.5 Competition and a shared purpose
In both biological and robotic collectives, competition can coexist with cooperation, as individuals pursuing local goals may still contribute to global functionality. Such emergent coordination depends not only on feedback regulation but on a shared purpose. At the microscopic scale, gene regulatory networks couple local signals to coordinated multicellular responses during processes like morphogenesis, embodying distributed feedback Grada et al. (2017); Angelini et al. (2011); Bi et al. (2016). At the macroscopic scale, local interaction rules (e.g., alignment, attraction, repulsion) are goal-conditioned: animals modulate the way they weigh neighbors and environmental cues in service of collective functions such as predator evasion, energy saving, or foraging Couzin et al. (2005); Portugal et al. (2014); Reebs (2000); Swaney et al. (2001). Robustness in living systems thus arises from distributed decision-making with a purpose, real-time adaptability driven by local feedback tuned to a common objective, rather than pre-planned, centralized control Marshall and Reina (2024).
Translating this principle to robotics suggests that achieving efficient, self-organizing collective motion requires embedding shared purpose directly into swarm control architectures. Incorporating goal-oriented feedback and distributed decision-making into physical design rules could enable robotic swarms to exhibit the same adaptability, resilience, and emergent coordination that characterize living systems.
III.6 Testing robot swarms
To test and refine swarm technologies, small-scale and inexpensive testbeds are crucialMokhtarian et al. (2025). As noted by M. Kegeleirs et al.Kegeleirs and Birattari (2025), one of the major obstacles to real-world swarm deployment is the lack of modern, affordable, and reliable experimental platforms. A number of swarm-specific testbeds have been developed, including E-Puck Mondada et al. (2009), Kilobot Rubenstein et al. (2012), Duckietown Paull et al. (2017), and Crazyflie Giernacki et al. (2017). These systems provide standardized, cost-effective environments for studying collective behavior at small scales.
Nevertheless, a common alternative is to build an in-house platform with custom-built García-Pérez et al. (2023) or commercial robots Mañas Álvarez et al. (2023). Among these, the Georgia Tech Robotarium Wilson et al. (2020) stands out as one of the most advanced and influential large-scale experimental facilities. Figure 2 shows a group of five robots. Each robot has an Arduino processor that controls the speed of the motors, receiving feedback signals from the wheel encoders. The Arduino communicates with a Raspberry Pi via a serial port. This Raspberry Pi receives the position or orientation of the robot relative to the arena robot from a server, thanks to an ArUCo-type marker. Some of the robots have battery sensors, and others have infrared distance sensors.
Figure 2: Group of five differential robots from the Robotarium-UCM.
Finally, efficient energy and resource management are critical for the long-term operation of swarm systems. Developing strategies to monitor and balance power consumption across agents will be essential for scaling swarm robotics toward real-world, autonomous applications.
IV Active Matter as an effective framework for Robophysics
Intelligent Soft Matter has been recently developed as a field of research aimed at creating biological-inspired materials with life-like capabilities, such as perception, learning, memory, and adaptive behaviorBaulin et al. (2025). Active matter has emerged in recent years as a robust framework for studying collective behaviors in both livingMarchetti et al. (2013); Bechinger et al. (2016); Gompper et al. (2025) and syntheticJu et al. (2025) systems. It focuses on ensembles of self-propelled agents, whether biological (e.g., bacteria, fish) or artificial (e.g., colloids), that consume energy to generate motion and stresses, thereby driving the system out of thermodynamic equilibrium Marchetti et al. (2013). These agents can exhibit striking emergent phenomena, such as flocking or swarming, that arise purely from local interactions. Synthetic micro/nanorobots could be have potential applications in biomedicine and environmental remediation.Ju et al. (2025) However, in contrast to biological systems, most active matter models lack internal decision-making and goal-oriented dynamics: while they capture emergent motion, they rarely encode the shared purposes that living collectives achieve, such as predator evasion, energy optimization, or efficient foraging Te Vrugt and Wittkowski (2025).
Recent developments highlight ongoing efforts to close this gap. New models seek to incorporate perception, adaptability, and purposeful behavior into synthetic active systems March-Pons et al. (2025, 2024); Acosta-Soba et al. (2025); Abdoli et al. (2025); Löwen and Liebchen (2025); Cai et al. (2025). At the same time, machine learning methods have been nowadays increasingly applied to active matter systemsBapst et al. (2020); Boattini et al. (2021); Clegg (2021); Wetzel et al. (2025); Jung et al. (2025); Janzen et al. (2024); Cichos et al. (2020); Janzen et al. (2023); Volpe et al. (2024). In particular, neural networks have been integrated into numerical simulations and trained on experimental data, enabling the simulations to capture complex, data-driven cell behaviors Minartz et al. (2025); neural network classifiers have been trained on numerical simulation data to distinguish motile from non-motile cells Braat et al. (2025); and graph neural networks have been used to extract the underlying pairwise interaction rules governing the dynamics of self-propelled colloids from experimental data of active colloidsRuiz-Garcia et al. (2024).
Moreover, beyond analysis, machine learning methods have been applied for inverse designLiu et al. (2025), which seeks to identify microscopic ingredients or control parameters leading to a desired macroscopic response. Reinforcement learningSutton et al. (1998); Brunton and Kutz (2022) (RL) is particularly well suited for this task, as it mimics trial-and-error strategies typical of biological systems. In Reinforcement Learning, an agent interacts with its environment, receiving positive or negative feedback that guides it toward actions maximizing a long-term reward Brunton and Kutz (2022). Such approaches have already been applied to soft and active matter systems, for example, to discover optimal protocols for steering active particles between states at minimal energetic cost Falk et al. (2021), to dynamically control the self-assembly of quasicrystalline structures Lieu and Yoshinaga (2025), and to learn control strategies for non-equilibrium dynamics using only local, imperfect gradient information Floyd et al. (2024).
V Outlook
Looking forward, we believe that active matter physics can be considered as the framework to tackle Robophysics. Active Matter allows us to extract the minimal ingredients needed to reproduce a Nature-inspired collective motion, such as purpose, perception, communication, and adaptive feedback, and to translate them into robotic swarms. Since machine learning and reinforcement learning have recently proven to be powerful tools for analyzing and controlling complex dynamical systems, their integration with active matter offers a path toward a conceptual framework that identifies the essential ingredients of living collectives and enables their implementation in robotic systems. Achieving this goal will require an interdisciplinary approach: physics to identify the governing principles, robotics and engineering to implement them in hardware, and computer science to provide the communication and data-processing infrastructure. By merging these perspectives, robophysics will move beyond locomotion toward programmable collectives that approach the functionality of living systems.
VI Acknowledgments
C.V. acknowledges fundings IHRC22/00002 and Proyecto PID2022-140407NB-C21 funded by MCIN/AEI /10.13039/501100011033 and FEDER, UE. DMF and GJ acknowledges support from MINECO (MCIN/AEI Grant No. PID2023-148991NA-I00). DMF also is funded through the Ramon y Cajal Program of the Spanish Ministry of Science and Innovation (Grant No. RYC2023-043002-I). J.J. and L.G-P. acknowledge support from iRoboCity2030-CM, Robotica Inteligente (TEC2024/TEC62TEC-2024/TEC-62), funded by the Programas de Actividades $I+D$ en Tecnologias en la Comunidad de Madrid. G.M acknowledges fundings Sapienza Award Horizon Europe - 2024.